LECT/PT 15-2018/SPRINGIADP Week 4: Assets and Statement Analysis / Chapter 11 Quiz Quiz 60 Brielle Phan, Attorney at Law purchased an office building with the accompanying land for $1,100,000. The latest tax assessment records value this property as follows: Land at $429,306 and Building at $579,214. How will Brielle allocate the cost of this property on the books of her law firm? Round percents to two decimal places, and final answer to the nearest dollar Cost allocated to Land: $ Cost allocated to Building: oe?? FS 8 5 6 8 9
media%2Fcca%2Fcca2598e-b9c1-4c9b-a160-62
media%2F46f%2F46fbed49-b46c-4f5b-844e-4a
Answer #1

Allocation of office building cost :

Land = 1100000*429306/1008520 = 468247

Building = 1100000*579214/1008520 = 631753

2) Straight line depreciation expense = (12200-1200)/5 =2200 per year

2011 dep = 2200

2012 dep = 2200

Accumlated dep at the end of 2012 = 2200*2 = 4400

Book value at the end of 2012 = 12200-4400 = 7800